solve: (x + y) (dx – dy) = dx + dy
_________________________________

Grade:12

1 Answers

Faiz
7 years ago
The expression can be written as:(x + y) / 1 = (dx + dy) / (dx - dy)Using componendo and dividendo:(x + y + 1) / (x + y - 1) = dx/dydy/dx = (x + y - 1) / (x + y + 1)Let x + y = t1 + dy/dx = dt/dx(t - 1) / (t + 1) + 1 = dt/dxdt/dx = 2t / (t + 1)(t + 1) / 2t dt = dxhere variables are separated and can be easily integrated to get:t/2 + 1/2 * ln|t| = x + c where t = x + y